OpenStudy (anonymous):

Determine the zeros of f(x) = x^3 – 12x^2 + 28x – 9.

OpenStudy (wasiqss):

find first zero by trial and error method

OpenStudy (anonymous):

x - 9 could be a factot or x - 3 using factor theorem f(9) = 9^3 - 12(81) + 28(9) - 9 using my calculator: this = 0 so (x-9) is a factor and 9 is a zero

OpenStudy (anonymous):

to find the others do long division or synthetic division

OpenStudy (wasiqss):

yes exactly

OpenStudy (anonymous):

or u can show me cus explaining thisgs to me doesnt help

OpenStudy (anonymous):

long division x^2 - 3x + 1 ----------------------- x - 9 ) x^3 - 12x^2 + 28x - 9 x^3 - 9x^2 ------------ - 3 x^2 + 28x - 3x^2 + 27x ------------- x - 9 x - 9 quotient is x^2 - 3x + 1 solve this to get the other zeros

OpenStudy (anonymous):

this wont factor so you need to use the quadratic formula